OpenStudy (anonymous):

multiply give your answer in lowes terms 14/25*15/44*8/21

OpenStudy (anonymous):

the numerator: 14*15*8 the denom.: 25*44*21 get those answers and then divide num/den.

OpenStudy (anonymous):

\[\frac{ 14 }{ 25 }*\frac{ 15 }{44 }*\frac{ 8 }{ 21 }=\frac{ 14*15*8 }{ 25*44*21 }\] \[\frac{ 14*15*8 }{ 25*44*21 }=\frac{ 1680 }{ 23100 }\]

OpenStudy (anonymous):

\[\frac{ 1680/420 }{ 23100/420 }=\frac{ 4 }{ 55 }\]

OpenStudy (anonymous):

yup